OpenStudy (amtran_bus):

Help me find complex zeros of x^2+2x+2

OpenStudy (amtran_bus):

I'm sure there are different methods, but can I use the quadratic formula?

OpenStudy (anonymous):

yes. or completing the square

OpenStudy (amtran_bus):

Can we work it both ways so I can be comfortable?

OpenStudy (amtran_bus):

That is where I am at for q formula.

OpenStudy (amtran_bus):

I figured it out on my own! YAY!

OpenStudy (anonymous):

ya your polynomial is ax^2 + bx + c quadratic: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } = \frac{ -2 \pm \sqrt{4 - 8} }{ 2 } = -1 \pm i\] completing square: consider only x^2 + bx: \[x^2 + bx = (x+\frac{ b }{ 2 })^2 - \left( \frac{ b }{ 2 } \right)^2\] \[x^2 + 2x + 2 =\left[ (x+1)^2 - 1 \right] + 2 = (x+1)^2 + 1\] finding roots of: (x+1)^2 + 1 = 0: \[(x+1)^2 = -1\]\[x+1 = \pm \sqrt{-1}\]\[x = -1 \pm \sqrt{-1} = -1 \pm i\] i just saw your quadratic. looks good. do you know how to simplify it into what i got?

OpenStudy (anonymous):

looks good. u got it. ignore last part lol. didnt scroll back before typing

OpenStudy (amtran_bus):

THANK YOU!

OpenStudy (amtran_bus):

Thanks for all you did, typing that out is a lot of work! I used my iPad and air server. lol

OpenStudy (anonymous):

hehe cool :) np